Really need help with this, thanks. 2 decimals

"Caspian Sea Drinks needs to raise $86.00 million by issuing additional shares of stock. If the market estimates CSD will pay a dividend of $1.48 next year, which will grow at 3.20% forever and the cost of equity to be 14.50%, then how many shares of stock must CSD sell?"
Answer #1
Price = Dividend in 1 year/(cost of equity - growth rate)
Price = 1.48/ (0.145 - 0.032)
Price = 13.0973451327

NUmber of shares = amount to raise/price = 86000000/13.0973451327 = 6566216